A TV special says there is a 10% incidence of sexually transmitted disease(STD)among all U.S. teens. A reporter fo the Rockville Bugle, looking for a story, surveyed 260 randomly chosen teenagers, and found that 39 had been to a clinic for STD treatment. Using a right tail test at a significance level of 0.05, the test statistic is:
A. 2.687
B. 2.758
C. 0.0256
D. 2.258

Answer by stanbon(75887)

A TV special says there is a 10% incidence of sexually transmitted disease(STD)among all U.S. teens. A reporter fo the Rockville Bugle, looking for a story, surveyed 260 randomly chosen teenagers, and found that 39 had been to a clinic for STD treatment. Using a right tail test at a significance level of 0.05, the test statistic is:
t(39/260) = (0.15-0.10)/[sqrt(0.10*0.90/260)] = 2.687
--------------
Cheeers,
Stan H.
--------------
A. 2.687
B. 2.758
C. 0.0256
D. 2.258